the rider rested after traveling 330 miles in 6 hours.if he traveled at the same rate the next day,how far could he go in 9 hours
Answer by jim_thompson5910(35256)

d = rt

330 = r*6

330/6 = r

55 = r

r = 55

-------------------------------------------------------

d = rt

d = 55*9

d = 495

So he can go 495 miles in 9 hours at the same rate.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------